# Ultra precise determination of Cs($nS_{1/2}$) and Cs($nD_{J}$) quantum defects for sensing and computing: Evaluation of core contributions

### v1

@@ Abstract
Neutral-atom-based quantum simulators and processors are the leading architectures for high fidelity quantum operations while atoms are being used extensively as sensors because of their stability and uniformity. Rydberg atoms are particularly suitable for quantum simulations and metrology, because their properties can be tuned through state selection and the use of electromagnetic fields. Precise knowledge of atomic structure enables engineering of atomic properties. Similarly for basic research, precise measurements of atomic energy levels provide stringent tests of atomic structure calculations, including quantum defect theories, and are critical for the determination of a number of fundamental constants in physics [1, 2, 3]. Cesium atoms have been used for searches of the electric-dipole moment of the electron [4, 5, 6, 7], amongst other fundamental quantities. Rapid developments in quantum measurement technologies, such as optical clocks [8, 9, 10, 11, 12], neutral-atom-based quantum computing [13, 14, 15, 16, 17] and Rydberg-atom-based electric field sensors [18, 19, 20] are adding impetus to accurately determine energy levels as well as derived quantities such as electric field polarizabilities and reduced electric-dipole matrix elements. The aim of this work is to improve the accuracy of Cs atomic energy level measurements for advanced applications in quantum sensing.
@@
The non-linear least squares fit to measured energies are shown in Fig. 1 with residues in the inset. The systematic downward trend for the higher principal quantum numbers of the $nD_{J}$ series is likely caused by DC Stark shifts. The quantum defects for all three series are presented in Table 1. Our results show that the $\delta_{2k}^{(l,J)}$ parameters for $k>2$ for all three series are insignificant within the measurement uncertainties. Several studies have reported these terms for the same states as non-zero at lower levels of experimental precision [45, 46]. Since the precision of our measurements is two orders of magnitude better than that of the prior work, it is likely that these terms have no physical significance and were fitting parameters that accommodated experimental uncertainties.
@@
We studied the fine-structure splitting of the $nD_{J}$ series. The fine-structure splitting, $\Delta_{\rm fs}$, can be expanded in terms of a power series in the effective principal quantum number, [$n-\tilde{\delta}^{(l,J=l\pm 1)}$], [53]:
$\Delta_{\rm fs}=\sum_{q=1}^{\infty}\frac{\xi^{(l)}_{q}}{[n-\tilde{\delta}^{(l,J=l\pm 1)}(n)]^{2q+1}}.$ (3)
The weighted average quantum defect, $\tilde{\delta}^{(l,J=l\pm 1)}(n)$, has the form
$\tilde{\delta}^{(l,J=l\pm 1)}(n)=w^{(l,J<l)}{\delta}^{(l,J<l)}(n)+w^{(l,J>l)}{\delta}^{(l,J>l)}(n).$ (4)
@@
Figure 3: The fine-structure splitting for $nD_{J}$. The red dashed line shows the prediction using Eq. (3). The fitting residuals are given as the black circles in the inset. The fine-structure interval based on the fitting coefficients reported in Ref. [53] is plotted as the blue dashed line for comparison. The corresponding residuals in data in Ref. [53], are shown as grey squares in the inset.
@@
The use of Eq. (3) to describe $\Delta_{\rm fs}$ had been proposed and used in previous work [53]. The physical meaning of $\xi^{(l)}_{q}$, have not been discussed to our knowledge. Here, we justify the expansion (3) by analytically deriving the spin–orbit coupling constant with hydrogenic wave functions, $\langle\frac{1}{r}\frac{{\rm d}V(r)}{{\rm d}r}{\textbf{L}}\cdot{\textbf{S}}\rangle$, for the atomic potential in the form of the sum of the Coulomb and polarization potentials, see Ref. [24]. Moreover, using the effective non-local model potential from Ref. [51] and the precise quantum defects, accurate wave functions and spin–orbit energies for the $nD_{5/2}$ and $nD_{3/2}$ series are determined. The numerical predictions $\Delta_{\rm fs}$ are obtained by taking the differences between the expectation values of the spin–orbit coupling operator of the two series [24]. They are fit to Eq. (3) to extract $\xi^{(l)}_{q}$, as shown in Table 3. The difference in the sign of $\xi^{(l)}_{2}$ may result from the model potential adopted for the numerical calculations. Importantly, the values of the leading coefficient for both the experimental and numerical approaches, $\xi^{(l)}_{1}$, are of similar order of magnitude. We note that the weighted average quantum defect, $\tilde{\delta}^{(l,J=l\pm 1)}(n)$, should be weighted by the total angular momenta of the $nD_{5/2}$ and $nD_{3/2}$ series, rather than being weighted equally as was done in Ref. [53].

### v2

@@ Abstract
Neutral-atom-based quantum simulators and processors are the leading architectures for high fidelity quantum operations while atoms are being used extensively as sensors because of their stability and uniformity. Rydberg atoms are particularly suitable for quantum simulations and metrology, because their properties can be tuned through state selection and the use of electromagnetic fields. Precise knowledge of atomic structure enables engineering of atomic properties. Similarly for basic research, precise measurements of atomic energy levels provide stringent tests of atomic structure calculations, including quantum defect theories, and are critical for the determination of a number of fundamental constants in physics [1, 2, 3]. Cesium atoms have been used for searches of the electric-dipole moment of the electron [4, 5, 6, 7], among other fundamental quantities. Rapid developments in quantum measurement technologies, such as optical clocks [8, 9, 10, 11, 12], neutral-atom-based quantum computing [13, 14, 15, 16, 17] and Rydberg-atom-based electric field sensors [18, 19, 20] are adding impetus to accurately determine energy levels as well as derived quantities such as electric field polarizabilities and reduced electric-dipole matrix elements. The aim of this work is to improve the accuracy of Cs atomic energy level measurements for advanced applications in quantum sensing.
@@
The non-linear least squares fit to measured energies are shown in Fig. 1 with residues in the inset. The systematic downward trend for the higher principal quantum numbers of the $nD_{J}$ series is likely caused by DC Stark shifts. The quantum defects for all three series are presented in Table 1. Our results show that the $\delta_{2k}^{(l,J)}$ parameters for $k>2$ for all three series are insignificant within the measurement uncertainties. Several studies have reported these terms for the same states as non-zero at lower levels of experimental precision [45, 46]. Since the precision of our measurements is two orders of magnitude better than that of the prior work, it is likely that these terms have no physical significance and were fitting parameters that accommodated experimental uncertainties [24].
@@
We studied the fine-structure splitting of the $nD_{J}$ series. The fine-structure splitting, $\Delta_{\rm fs}$, can be expanded in terms of a power series in the effective principal quantum number, [$n-\tilde{\delta}^{(l,J=l\pm 1/2)}$], [53]:
$\Delta_{\rm fs}=\sum_{q=1}^{\infty}\frac{\xi^{(l)}_{q}}{[n-\tilde{\delta}^{(l,J=l\pm 1/2)}(n)]^{2q+1}}.$ (3)
The weighted average quantum defect, $\tilde{\delta}^{(l,J=l\pm 1/2)}(n)$, has the form
$\tilde{\delta}^{(l,J=l\pm 1/2)}(n)=w^{(l,J<l)}{\delta}^{(l,J<l)}(n)+w^{(l,J>l)}{\delta}^{(l,J>l)}(n).$ (4)
@@
Figure 3: The fine-structure splitting for $nD_{J}$. The red dashed line shows the prediction using Eq. (3). The fitting residuals are given as the black circles in the inset. The fine-structure interval based on the fitting coefficients reported in Ref. [53] is plotted as the blue dashed line for comparison. The corresponding residuals in data in Ref. [53], are shown as gray squares in the inset.
@@
The use of Eq. (3) to describe $\Delta_{\rm fs}$ had been proposed and used in previous work [53]. The physical meaning of $\xi^{(l)}_{q}$, have not been discussed to our knowledge. Here, we justify the expansion (3) by analytically deriving the spin–orbit coupling constant with hydrogenic wave functions, $\langle\frac{1}{r}\frac{{\rm d}V(r)}{{\rm d}r}{\textbf{L}}\cdot{\textbf{S}}\rangle$, for the atomic potential in the form of the sum of the Coulomb and polarization potentials, see Ref. [24]. Moreover, using the effective non-local model potential from Ref. [51] and the precise quantum defects, accurate wave functions and spin–orbit energies for the $nD_{5/2}$ and $nD_{3/2}$ series are determined. The numerical predictions $\Delta_{\rm fs}$ are obtained by taking the differences between the expectation values of the spin–orbit coupling operator of the two series [24]. They are fit to Eq. (3) to extract $\xi^{(l)}_{q}$, as shown in Table 3. The difference in the sign of $\xi^{(l)}_{2}$ may result from the model potential adopted for the numerical calculations. Importantly, the values of the leading coefficient for both the experimental and numerical approaches, $\xi^{(l)}_{1}$, are of similar order of magnitude. We note that the weighted average quantum defect, $\tilde{\delta}^{(l,J=l\pm 1/2)}(n)$, should be weighted by the total angular momenta of the $nD_{5/2}$ and $nD_{3/2}$ series, rather than being weighted equally as was done in Ref. [53].